NEXT
Make New Entry,
Make Followup Entry
User name E.Chudakov
Log entry time 01:00:22 on November 26, 2008
Entry number 250787
keyword= Experiment summary for Nov 24(swing) - 25(swing)
Summary for Nov 24 (swing) - Nov 25 (swing)
Smooth data for Nov 24 swing and Nov 25 owl. The target cell changed on
Nov 25 day. Beam is back at about 21:40. A regular Compton tune performed
(without any tilt as before).
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/11/24 swing | 5.95 |
0.32 | 1.73 |
2008/11/25 owl | 7.08 |
0.00 | 0.92 |
2008/11/25 day | 0.28 |
3.09 | 4.62 |
2008/11/25 swing | 0.58 |
5.66 | 1.76 |
total | 13.89 |
9.07 | 9.03 |
Main events/issues:
- The regular Compton tune, without a previous tilt, was successful.
The background in the photon detector from the Hall B beam is
still high.
Closing the Hall A slit and reducing the attenuation reduced the
background by a factor of 20.
A good Compton spectrum from the photodetector is demonstrated.
The source of noise in the electronic detector is found.
Plans for the next few days:
- Measurements with the reference cell, outlined on the white board.
- Production.